# QUIJOTE-TFGI polarization calibration Ground characterization and on-sky validation with Tau A and the Moon

### v1

@@ 4.1.2 Responsivity calibration with Tau A
where $\Omega_{b}$ is the solid angle subtended by the model beam, $\Omega_{b}=4.77\times 10^{-5}\penalty 10000\ \mathrm{sr}$. The expected temperature at 31 GHz is then 230 mK. We obtain the responsivity by dividing the amplitude calculated through BF photometry on the intensity map, $I$, by the calculated temperature. We report the responsivities derived from Tau A in Tables 7 (for the single raster scan described in the previous section). We also calculated the responsivity values for the coadded monthly maps (from Fernández-Torreiro, M. et al. in prep.) from November 2021, December 2021, and January 2022. In this way, we have a sky-sky comparison between the responsivities calculated from Tau A and the Moon, as we will discuss in Sect. 5.2. We report the responsivities derived from Tau A in Tables 7 (for the single raster scan described in the previous section) and 10 (for the monthly coadded maps).
@@ 4.1.3 Polarization angle
We obtain estimates of $Q$ and $U$ ($Q_{\rm BF}$ and $U_{\rm BF}$) from the BF analysis described in Sect. 4.1.1. The results obtained after introducing these values into Eq. (10) are summarized in Table 6. Finally, we estimated the uncertainties of $\phi_{\rm{c}}$ by drawing 10 000 random realizations for $Q$ and $U$. The width of the distribution of the $\phi_{\rm{c}}$ values from these realizations was utilized to estimate the uncertainty of this parameter.
@@ 4.2 The Moon
We use three Moon observations to characterize TGI Pixel 23 in the sky. Each observation involves repeated azimuth slews (azimuth scans at fixed elevation) lasting $\sim$1 hour, covering an azimuth range of 100 deg at a speed of 4.5 deg/s.
@@ 4.2.1 Moon emission model
where $T_{0}$ represents the mean temperature and $T_{1}$ is the amplitude of the temperature variation associated with the Moon’s phase, $\phi(t)$. Finally, $\xi(\nu)$ is the phase relative to the time of the full Moon. The values at 31 GHz for these parameters are interpolated from the values presented in Hafez et al. (2014), obtaining: $T_{0}(31\,\rm{GHz})=213.3\pm 0.9$ K, $T_{1}(31\,\rm{GHz})=35\pm 2$ K $\xi(31\,\rm{GHz})=28\pm 4$ deg.
@@
Figure 7: Schema describing the Moon’s emission model and the parameters involved: the refraction index of the Moon ($n_{\rm{Moon}}$), the refraction index of free space ($n_{t}$), the incident ($\theta_{i}$) and transmitted ($\theta_{t}$) angles, the radial coordinate ($r^{\prime}$), and the radius of the Moon ($R$).
By exploiting the transmittance values described in Eqs. (12), we calculate the intensity ($I_{\rm{Moon}}$) and the polarization ($P_{\rm{Moon}}$) of the Moon’s surface as a function of its temperature ($T_{\mathrm{Moon}}$), assuming it is uniform across the entire surface, as (Bischoff 2010):
@@
where the term $\partial\theta_{i}/\partial\theta_{j}$ describes the change in the solid angle of the incident beam after the refraction process.
@@
As in Xu et al. (2020), we do not observe the emergence of the quadrupolar polarization pattern in a single observation of the Moon for the central detector. In this work, we assume a uniform temperature, $T_{\rm{Moon}}$, across the lunar disk, thereby simplifying the latitude-dependent model proposed by Zheng et al. (2012). Even after including the latitude-dependent model and stacking the three available maps taken at different epochs and incorporating the latitude-dependent temperature model, the quadrupolar pattern is not recovered in Pixel 23 (central pixel). The behavior is repeatable and stable over the timescale of our observations ($\sim$2 months). In contrast, the quadrupolar structure is clearly visible in single observations from other TFGI detectors. We performed preliminary simulations and tests to investigate potential explanations, including intensity-to-polarization leakage, non-linear response, saturation effects, phase state angle misestimation, and mapping strategy artifacts. None of these effects explains the observed behavior. Therefore, the origin of this discrepancy for the central detector remains unclear. A dedicated future study will address this issue in detail, including a comprehensive analysis of Moon observations across all TFGI detectors, once the instrument is installed again on the focal plane of QT2.
Nevertheless, it is important to stress that the expected rotations between the polarization patterns observed for Stokes $Q$ and $U$, as well as across different channels, are maintained even if the quadrupolar pattern is not visible (it misses one lobe). Thus, the measurements are still providing a robust determination of the polarization angle.
@@ 4.2.4 Polarization angle
$\chi^{2}(\phi_{\rm{c}})=$ $\frac{1}{\sigma_{\rm{DET}}^{2}}\sum_{p}\bigg[\left(\dfrac{Q_{\rm obs,\textit{p}}}{\max(I^{\prime})}-\dfrac{Q_{p}(2\phi_{\rm{c}})}{\max(I)}\right)^{2}+$ (16)
$+\left(\dfrac{U_{\rm obs,\textit{p}}}{\max(I^{\prime})}-\dfrac{U_{p}(2\phi_{\rm{c}})}{\max(I)}\right)^{2}\bigg].$ (17)
where the sum runs over all pixels $p$ inside an aperture of radius 1 deg, $\sigma_{\rm DET}$ is is the detector’s noise, $\max(I)$ and $\max(I^{\prime})$ are the maximum intensity values measured on the respective maps and $Q_{p}(2\phi_{\rm{c}})$ and $U_{p}(2\phi_{\rm{c}})$ are the $2\phi_{\rm{c}}$ rotations of the modeled maps:
@@
Finally, we note that the resulting fitting error is small, with a typical uncertainty of $0.04$ deg. However, this estimate does not include systematic effects, which are expected to dominate the total error budget. Overall, the methodology appears promising; with further development, it could become a precise tool for characterizing the polarization angle in ground-based experiments.
@@ 5.2 Responsivity
Figure 9 presents the results on responsivity as a function of the observation period for each channel, including Tau A and the Moon. When comparing Tau A and the Moon, the responsivity is well in accordance. Overall, intercalibration within the channels is maintained, but the responsivity shows significant variations over time. Furthermore, the results are coherent within the period and astronomical source, suggesting a common factor that alters the gain, such as the temperature of the backend.
@@ 6 Conclusions
In particular, the Moon observations provide an independent check on the instrumental polarization calibration. By assuming a fixed polarization efficiency from Tau A, we estimate the effective refraction index of the Moon at 31 GHz to be $n_{\rm{Moon}}=1.209\pm 0.007\,\text{(stat)}\,\pm 0.005\,\text{(sys)}$, consistent with limb-weighted emission dominated by uppermost surface layer and previous measurement. This result aligns with previous findings that the apparent dielectric constant can vary significantly with surface roughness, reinforcing the Moon’s utility as a polarization calibrator at microwave frequencies.
This work highlights the need for implementing a real-time system to monitor responsivity in conjunction with calibrator observations. Following the example of the MFI, we will install an internal diode calibrator to sample responsivity every 30 seconds. This system will provide relative responsivity calibration for tracking stability, while absolute calibration will continue to rely on astrophysical sources, such as Tau A. The new diode system will enhance the mitigation of systematics, yielding more robust polarization angle determination and improved temporal stability of the instrument’s response.
@@ Appendix A Effective refraction index of the Moon at 31 GHz under smooth-surface assumption
From this procedure, we derive an effective refraction index of $n_{\rm{Moon}}=1.209\pm 0.007\,\text{(stat)}\,\pm 0.005\,\text{(sys)}$. To estimate the systematic uncertainty (sys) for this measurement, alongside the statistical uncertainty (stat), we rerun the calculation varying the beam FWHM by 2%, based on the ellipticity beam calculation of TGI (see Fernández-Torreiro 2023). The total uncertainty on the refraction index is $\sim$0.01. Additional factors, such as the assumed polarization model, viewing geometry, surface heterogeneity, and limb dominance, which are not included in this analysis, may also contribute to systematic uncertainties.

### v2

@@ 4.1.2 Responsivity calibration with Tau A
where $\Omega_{b}$ is the solid angle subtended by the model beam, $\Omega_{b}=4.77\times 10^{-5}\penalty 10000\ \mathrm{sr}$. The expected temperature at 31 GHz is then 230 mK. We obtain the responsivity by dividing the amplitude calculated through BF photometry on the intensity map, $I_{\rm BF}$, by the calculated temperature. We report the responsivities derived from Tau A in Table 4 (for the single raster scan described in the previous section). We also calculated the responsivity values for the coadded monthly maps (from Fernández-Torreiro, M. et al. in prep.) from November 2021, December 2021, and January 2022. We report the responsivities derived from Tau A in Tables 4 (for the single raster scan described in the previous section) and 10 (for the monthly coadded maps). We also compute the responsivity from Moon observations in the following Section; we present this sky-sky check of the responsivity in Sect. 5.2.
@@ 4.1.3 Polarization angle
We compute $\phi_{c}$ from $Q_{\rm BF}$ and $U_{\rm BF}$ values from Sect. 4.1.1 using Eq. (10). We summarize the results in Table 6. Finally, we estimate the uncertainties of $\phi_{\rm{c}}$ by drawing 10 000 random realizations for $Q$ and $U$ distributions. We define these distributions as Gaussians with centers equal to $Q_{\rm BF}$ and $U_{\rm BF}$ and standard deviations equal to $\sigma(Q_{\rm BF})$ and $\sigma(U_{\rm BF})$. We evaluated $\phi_{\rm c}$ for each pair $Q_{\rm BF}-U_{\rm BF}$ and then built its distribution. We defined the uncertainty on $\phi_{\rm c}$ as the width of such a distribution.
@@ 4.2 The Moon
We use three Moon observations to characterize TGI Pixel 23 in the sky. These observations were performed on 21 November 2021, 28 December 2021, and 23 January 2022. Each observation involves repeated azimuth slews (azimuth scans at fixed elevation) lasting $\sim$1 hour, covering an azimuth range of 100 deg at a speed of 4.5 deg/s. The azimuth scan spans 100 deg to ensure full beam crossing for all pixels, thereby minimizing edge effects and ensuring uniform beam sampling.
@@ 4.2.1 Moon emission model
where $T_{0}$ represents the mean temperature and $T_{1}$ is the amplitude of the temperature variation associated with the Moon’s phase, $\phi(t)$. Finally, $\xi(\nu)$ is the phase relative to the time of the full Moon. The values at 31 GHz for these parameters are interpolated from the values presented in Hafez et al. (2014), obtaining: $T_{0}(31\,\rm{GHz})=213.3\pm 0.9$ K, $T_{1}(31\,\rm{GHz})=35\pm 2$ K $\xi(31\,\rm{GHz})=28\pm 4$ deg. For our three observing dates (21 November 2021, 28 December 2021, and 23 January 2022), the corresponding lunar phases are $\phi$=19.4 deg, 104.4 deg, and 60.2 deg, respectively. In this work, we do not include spatial dependence on the thermal signal due to the low angular resolution of the telescope, and it is not expected to significantly affect either the absolute signal (at most $\sim$8% discrepancy with Zheng et al. 2012 model in these dates) or the polarization morphology. Additionally, our observations were made relatively near the time of the full Moon, which also decreases the contribution of spatial anisotropies.
@@
Figure 7: Schematic describing the Moon’s emission model and the parameters involved: the refraction index of the Moon ($n_{\rm{Moon}}$), the refraction index of free space ($n_{t}$), the incident ($\theta_{i}$) and transmitted ($\theta_{t}$) angles, the radial coordinate ($r^{\prime}$), and the radius of the Moon ($R$).
By exploiting the transmittance values described in Eqs. (12), we calculate the intensity ($I_{\rm{Moon}}$) and the polarization ($P_{\rm{Moon}}$) of the Moon’s surface as a function of its temperature ($T_{\mathrm{Moon}}$):
@@
where the term $\partial\theta_{i}/\partial\theta_{t}$ describes the change in the solid angle of the incident beam after the refraction process.
@@
As in Xu et al. (2020), we do not observe the emergence of the quadrupolar polarization pattern in a single observation of the Moon for the central detector. In this work, we assume a uniform temperature, $T_{\rm{Moon}}$, across the lunar disk, thereby simplifying the latitude-dependent model proposed by Zheng et al. (2012). Even after including the latitude-dependent model and stacking the three available maps taken at different epochs and incorporating the latitude-dependent temperature model, the quadrupolar pattern is not recovered in Pixel 23 (central pixel). The behavior is repeatable and stable over the timescale of our observations ($\sim$2 months). In contrast, the quadrupolar structure is clearly visible in single observations from other TFGI detectors. We performed preliminary simulations and tests to investigate potential explanations, including intensity-to-polarization leakage, non-linear response, saturation effects, phase state angle misestimation, and mapping strategy artifacts. None of these effects explains the observed behavior. Therefore, the origin of this discrepancy for the central detector remains unclear. Even when the quadrupolar pattern is only partially visible (i.e., one lobe is missing), the relative orientation between the polarization patterns observed in Stokes $Q$ and $U$ is preserved. Since the polarization angle determination relies on these relative rotations rather than on the full quadrupolar morphology, the measurements still provide a robust estimate of the polarization angle. A dedicated future study will address this issue in detail, including a comprehensive analysis of Moon observations across all TFGI detectors, once the instrument is installed again on the focal plane of QT2. The TFGI instrument was temporarily removed after the initial commissioning phase. A new configuration comprising 19 receivers (10 at 31 GHz and 9 at 41 GHz) is being prepared, and observations are expected to resume in spring 2026.
@@ 4.2.4 Polarization angle
$\chi^{2}(\phi_{\rm{c}})=$ $\frac{1}{\sigma_{\rm{DET}}^{2}}\sum_{p}\bigg[\left(\dfrac{Q_{\rm obs,\textit{p}}}{\max(I_{\rm obs})}-\dfrac{Q_{p}(2\phi_{\rm{c}})}{\max(I)}\right)^{2}+$ (16)
$+\left(\dfrac{U_{\rm obs,\textit{p}}}{\max(I_{\rm obs})}-\dfrac{U_{p}(2\phi_{\rm{c}})}{\max(I)}\right)^{2}\bigg].$
where the sum runs over all pixels $p$ inside an aperture of radius 1 deg, $\sigma_{\rm DET}$ is is the detector’s noise, $\max(I)$ and $\max(I_{\rm obs})$ are the maximum intensity values measured on the respective maps and $Q_{p}(2\phi_{\rm{c}})$ and $U_{p}(2\phi_{\rm{c}})$ are the $2\phi_{\rm{c}}$ rotations of the modeled maps:
@@
Finally, we note that the resulting fitting error is small, with a typical uncertainty of $0.04$ deg. However, this estimate does not include systematic effects, which are expected to dominate the total error budget. Model-related uncertainties, including those linked to the latitude-dependent temperature model, are not considered in this work. Preliminary simulations indicate that these effects do not significantly influence the results; however, a complete evaluation of such model uncertainties is beyond the scope of this work and will be addressed in future dedicated studies. Overall, the methodology appears promising; with further development, it could become a precise tool for characterizing the polarization angle in ground-based experiments.
@@ 5.2 Responsivity
Figure 9 presents the results on responsivity as a function of the observation period for each channel, including Tau A and the Moon. When comparing Tau A and the Moon, the responsivity presents discrepancies of $\sim$15%. Overall, intercalibration within the channels is maintained, but the responsivity shows significant variations over time. Furthermore, the results are coherent within the period and astronomical source, suggesting a common factor that alters the gain, such as the temperature of the backend.
@@ 6 Conclusions
In particular, the Moon observations provide an independent check on the instrumental polarization calibration. By assuming a fixed polarization efficiency from Tau A, we estimate the effective refraction index of the Moon at 31 GHz to be $n_{\rm{Moon}}=1.209\pm 0.007\,\text{(stat)}\,\pm 0.009\,\text{(sys)}$, consistent with limb-weighted emission dominated by uppermost surface layer and previous measurement. This result aligns with previous findings that the apparent dielectric constant can vary significantly with surface roughness, reinforcing the Moon’s utility as a polarization calibrator at microwave frequencies.
This work highlights the need for implementing a real-time system to monitor responsivity in conjunction with calibrator observations. This conclusion is motivated by the observed temporal variability of calibration parameters during commissioning, which demonstrates that continuous monitoring would be necessary to track and correct such effects in real time. Following the example of the MFI, we will install an internal diode calibrator to sample responsivity every 30 seconds. This system will provide relative responsivity calibration for tracking stability, while absolute calibration will continue to rely on astrophysical sources, such as Tau A. The new diode system will enhance the mitigation of systematics, yielding more robust polarization angle determination and improved temporal stability of the instrument’s response.
The scientific goal is to achieve an absolute calibration performance for TFGI comparable to that obtained with the QUIJOTE MFI 11 GHz band (Rubiño-Martín et al. 2023), namely an internal calibration uncertainty of $\sim$1% and an overall uncertainty of $\sim$5% when including systematic errors associated with the modeling of the calibrators. For the polarization angle, we aim for an accuracy of about 1 deg, similar to what is achieved at 11 GHz with MFI. In this work, the absolute responsivity (derived independently from Tau A and from the Moon) is consistent at the few-percent level. However, no internal calibration diode is currently installed, and therefore, the absolute responsivity is not yet stabilized across different observing periods. The internal consistency of the responsivity calibration is presently limited to <<8% level due to residual systematic effects identified during commissioning. Regarding the polarization angle, we obtain uncertainties at the $\sim$1 deg level, well within the <<3 deg requirement inferred from previous QUIJOTE analyses (López Caraballo 2013). These performances are adequate for the primary QUIJOTE science goals at these frequencies.
@@ Appendix A Effective refraction index of the Moon at 31 GHz under smooth-surface assumption
From this procedure, we derive an effective refraction index of $n_{\rm{Moon}}=1.209\pm 0.007\,\text{(stat)}\,\pm 0.009\,\text{(sys)}$ To estimate the systematic uncertainty, we first reran the calculation varying the beam FWHM by 2%, based on the TGI beam ellipticity analysis (see Fernández-Torreiro 2023), obtaining a contribution of 0.005. In addition, we considered potential intensity-to-polarization leakage. Assuming a conservative worst-case scenario of 0.5% leakage from intensity into both $Q$ and $U$, as expected during high-speed telescope scanning (Fernández-Torreiro 2023), this effect contributes an additional 0.008 to the refraction index. These two contributions are combined in quadrature, yielding the quoted systematic uncertainty of 0.009. Additional factors, such as the assumed polarization model, viewing geometry, surface heterogeneity, and limb dominance, which are not included in this analysis, may also contribute to systematic uncertainties. The quoted systematic uncertainty should therefore be regarded as a lower limit.